answersLogoWhite

0

🎒

Probability

The probability of a certain event is a number expressing the likelihood that a specific event will occur, expressed as the ratio of the number of actual occurrences to the number of possible occurrences. In mathematics, it is a measure of how often an event will happen and is the basis of statistics.

14,643 Questions

What is the probability of getting a diamond in a deck of 52 cards?

-- Total number of cards = 52.

-- Number of diamonds = 13 .

-- Probability of success = (# of successes)/(# of total possibilities) = 13/52 = 25% .

What is the probability of rolling a 5 on a 6 sided dice?

There is always a 1 in 6 chance of rolling ANY number on a six-sided dice, as there are 6 numbers.

What are the odds of rolling a 6 or an 8 before rolling a 7 in craps?

Here are the odds in craps: 7 vs. 6 Odds: 6 to 5 7 vs. 5 Odds: 6 to 4 or 3 to 2 7 vs. 4 Odds: 6 to 3 or 2 to 1 7 vs. 3 Odds: 6 to 2 or 3 to 1 7 vs. 2 Odds: 6 to 1

What are the odds of winning the Mega Millions Jackpot?

The odds of winning prizes are shown in table from the official Mega Millions website. From the table, you can calculate the cumulative probability of winning a prize and the expectation of the prize.

- If you bought 1 ticket

You have 1/75 + 1/141 + 1/844 + 1/306 + 1/13781 + 1/15313 + 1/689065 + 1/3904701 + 1/175711536 = 1/40 chances to win a prize with the expectation of $2/75 + $3/141 + $10/844 + $7/306 + $150/13781 + $150/15313 + $10,000/689065 + $250,000/3904701 + Jackpot/175711536 = $0.83. Here we assumed Jackpot prize is $12,000,000. This can be interpreted statistically as follows: You pay $1 to buy 1 ticket, and you have 2.5% of chances to get $0.83 back out of your $1 and 97.5% of chances to lose your $1. So, the expected net worth of the 1 ticket is 0.025x$0.83=$0.02, which is about 2% of the purchased ticket price.* In case the Jackpot is rolled over, the expectation of the prize will increase, but increasing the Jackpot prize from $12,000,000 to $120,000,000 only results in increasing the expectation of the prize from $0.83 to $1.44, thus the expected net worth of a ticket increases from 0.025x$0.83=$0.02 to 0.025x$1.44=$0.036, which is a trivial difference.

* The calculations are brought from PickSmarter.com.

- If you bought 2 tickets

The chances for at least one of 2 tickets to win a prize is 1-(39/40)2 =0.049 with the expectation of $1.65. So, you pay $2 to buy 2 tickets, and you have 4.9% of chances to get $1.65 back out of your $2 and 95.1% of chances to lose your $2. The expected net worth of the 2 tickets is 0.049x$1.65=$0.08, which is about 4% of the purchased tickets price.

- If you bought 40 tickets

The chances for at least one of 40 tickets to win a prize is 1-(39/40)40 =0.637 with the expectation of $33.04. So, you pay $40 to buy 40 tickets, and you have 63.7% of chances to get $33.04 back out of your $40 and 36.3% of chances to lose your $40. The expected net worth of the 40 tickets is 0.637x$33.04=$21.05, which is about 52% of the purchased tickets price.

- If you bought 100 tickets

The chances for at least one of 100 tickets to win a prize is 1-(39/40)100 =0.92 with the expectation of $83. So, you pay $100 to buy 100 tickets, and you have 92% of chances to get $83 back out of your $100 and 8% of chances to lose your $100. The expected net worth of the 100 tickets is 0.92x$83=$76.4, which is about 76% of the purchased tickets price.

So, the more tickets you buy, the chances of at least one of your tickets win a prize increase with higher expected amount of prize, and thus the chances of your getting nothing decrease, but the amount of money you lose increases when you do lose. The percentage of the expectected net worth of the ticksts gradually converges to the purchased tickets price, at the risk of losing more money with less probability.

How do you use probable in a sentence?

Probable is the same as likely or that something is likely to happen. While I hope it doesn't rain, the dark clouds in the sky make it probable that it will.

For a sample of 140 randomly selected patients the mean amount spent was 86.50 and the standard deviation was 11.45. What is a 95 percent confidence interval for the mean?

sample size, n = 140
standard deviation, s = 11.45

standard error of the mean, SE = s / n^1/2 = 11.45 / 140^1.2 = 0.9677

95% confidence interval => mean +- 1.96SE

95% CI = 86.5 - 1.96*0.9677; 86.5 + 1.96*0.9677

= 84.6; 88.4

How does PCIe Version 2 compare to Version 1?

PCIe version 2 doubled the fequency of the PCIe bus, theoretically doubling the throughput, It also allows for up to 32 lanes on one slot

When two dice are rolled find the probability of not rolling a sum of 5?

When rolling a pair of dice there are 6∙6 = 36 possible outcomes.

The outcomes that give a sum of 5 are 4, [(1,4), (2,3), (3,2), (4,1)].

So the probability of not rolling a sum of 5 is:

P(NOT 5) = 1 - P(5) = 1 - 4/36 = 32/36 = 8/9 = 0.8888... ≈ 88.9%

A coin is tossed twice If heads comes up either time you get 2 dollars If heads does not come up you lose 4 dollars What is the expected value of this game?

There is a .25 chance that no heads will come up. (1/2*1/2=.25)

Therefore, there is a .75 chance that one or more heads will come up.

Value of game = .75*$2-.25*$4=$1.5-$1=$.50

Over time, you should come out ahead.

What is the probability density function for a uniforn two-dimentional random variable?

Your question is a bit unclear. I hope this answers your question, but if not, ask it again in a more specific manner.

The pdf of a continuous uniform random variable (X) distributed over the range a to b is f(x) = 1/(b-a).

If we consider x and y are dimensions, and each have a uniform distribution, then we can similarly define f(y) = 1/(c - d). Now, since these are independent, the joint distribution is f(x,y) = 1/(c-d) * 1/(b-a).

So, if we want to find the probability of X < x* and Y < y*, where x* is in the range of a to b and y* is in the range of c to d, we would integrate twice over f(x,y):

F(x*,y*) = integral from a to x* , integral from c to y* f(x,y) dy dx

How many event in a decatholon?

There are ten events in a decathlon because deca- means ten

How many possible permutations using the word bite?

4!

Four factorial.

4 * 3 * 2

= 24 permutations

------------------------

What does responsiblity mean?

It is when you have to take on a huge role and not disapoint yourself. Say if you had to babysit a child you will need to have full consertration and need to be responsible if the child gets hurt.

How high of number dice sets are available?

The highest number on a single "die" from a dice set is six. Therefore, the highest number available that one could roll on a dice set would be 6 multiplied by the number of dice thrown.

What are the odds of getting a sum divisible by by 4 on dice?

there are 36 possible outcomes and 9 ways to get sums divible by 4, so odds are 1 in 4

Trending Questions
How many ways can the letter of the word probe be arranged? What is the probability of getting a 2 with one roll of the dice? In how many ways can 3 clubs be drawn from a standard deck of 52 cards? What are power pack heads? What is the probability of not rolling a 8 on dice? Why does the standard deviation get smaller as the individual in a group score more similarly on a test? What is the probability that one of the offspring mice will have black fur? What are the odds of selecting a four of diamonds out of a deck of fiffty-two card? Is defined as any adverse event that is likely to occur? Three coins are tossed simultaneously the probability of getting atmost one head? When 2 dice are rolled what is the probability of getting a sum less than 4 or greater than 9? A pizza has eight possible toppings. three of the possible toppings are meat and five are vegetarian. if two toppings are picked what is the probability that at least one of them is meat? For a large target population what are the sampling methods use to select the exact population? What's H aggregate? What is the probability of rolling a 3 and then 1 of six side dice? What are the odds of having twins with no family history? What is the probability of the spinner not landing on 1 or 2? What is the probability of getting heads one-half N times when flipping a coin N times? Squashed circle with no straight edges? How many possible outcomes are there when you roll two number cubes and toss one coin?